# Robust mass lumping and outlier removal strategies in isogeometric analysis

### v1

@@ 4.2 Eigenvalue and eigenvector computations
Algorithm 1: Lanczos method [55, Algorithm 9.1]
The vectors $\mathbf{v}_{j}$ computed during the course of the iterations are stored along the columns of the matrix $V_{m}$; i.e. $V_{m}=[\mathbf{v}_{1},\mathbf{v}_{2},\dots,\mathbf{v}_{m}]$, which by construction forms a $B$-orthonormal basis for the Krylov subspace
$\mathcal{K}_{m}(C,\mathbf{b})=\operatorname{\operatorname{span}}(\mathbf{b},C\mathbf{b},C^{2}\mathbf{b},\dots,C^{m-1}\mathbf{b}),$
where $C=B^{-1}A$ and $\mathbf{b}$ is the starting vector, usually randomly chosen. The coefficients $\alpha_{j}$ and $\beta_{j}$ are stored in the symmetric tridiagonal matrix
$T_{m}=\begin{pmatrix}\alpha_{1}&\beta_{2}&&\\ \beta_{2}&\alpha_{2}&\ddots&\\ &\ddots&\ddots&\beta_{m}\\ &&\beta_{m}&\alpha_{m}\end{pmatrix}.$
Rewriting line 5 in matrix form leads to the famous Lanczos decomposition
$AV_{m}=BV_{m}T_{m}+\beta_{m+1}B\mathbf{v}_{m+1}\mathbf{e}_{m}^{T}.$
From the $B$-orthonormality of $V_{m}$, we deduce that
$V_{m}^{T}AV_{m}=T_{m},\quad V_{m}^{T}BV_{m}=I_{m}.$
Following the Rayleigh-Ritz procedure [53, 55], an approximate eigenpair (called Ritz pair) is defined as $(\lambda,\mathbf{u})=(\mu,V_{m}\mathbf{q})$, where $(\mu,\mathbf{q})$ is an exact eigenpair of the much smaller tridiagonal matrix $T_{m}$.
@@
Although we have implemented an eigensolver from scratch for the sake of writing this paper, it is absolutely not necessary for applying the techniques presented herein. Several efficient implementations are available in software packages and all the user has to worry about is supplying an algorithm for computing matrix-vector multiplications with the stiffness matrix and solving linear systems with the (lumped) mass matrix, which depends on the nature of the problem (e.g. single-patch, multipatch,…). If the number of iterations $m$ remains relatively small, which is ensured through restarting procedures, each iteration of the Lanczos method costs nearly as much as an iteration of the central difference method. Moreover, the Lanczos method is known to converge very fast to eigenvalues that are well separated from the rest of the spectrum [59, 53, 55], which is precisely a distinctive feature of outliers. However, the time span of the simulation must be sufficiently large to amortize the cost of computing outlier eigenpairs. In general, the shorter the simulation, the smaller the number of outliers we can afford computing. Finally, although our method scales down outlier frequencies, it does not remove the corresponding outlier modes that might negatively impact the solution.
The deflation procedure proposed in this section is very general and can in principle be applied to any type of problem, including nontrivial (multipatch) geometries.
In the multipatch setting, it is possible to locally scale the single-patch system matrices, similarly to local (elementwise) mass scaling techniques [27, 60]. Indeed, Lemma 3.22 reveals that the largest eigenvalues of $(\mathcal{K},\mathcal{P}_{i})$ could be controlled by the largest eigenvalues of $(\mathcal{K}_{r},\mathcal{P}_{r,i})$ and suggests a scaling strategy directly targeting the origin of the issue: at the patch level. While this strategy is generally cheaper than scaling $(\mathcal{K},\mathcal{P}_{i})$ globally, it has three shortcomings: firstly, the assembly into global matrices generally effects the smallest eigenvalues; secondly it cannot remove outliers introduced by the $C^{0}$ coupling of patch interfaces and finally, since the inverse of the global mass matrix is not given by the assembly of the local inverses, we cannot directly apply (4.1) locally. To resolve the third issue, we suggest locally scaling the stiffness matrix instead. Denoting $\bar{\mathcal{K}}_{r}$ the locally scaled stiffness matrix and recalling that the perturbation is negative semidefinite (see Section 4.1), $\bar{\mathcal{K}}_{r}\preceq\mathcal{K}_{r}$ and
@@
For some special cases it is yet far more advantageous to exploit the structure of the problem. In particular, for problems featuring a Kronecker product structure, all outliers can be removed by separately scaling the 1D factor matrices. For maximally smooth spline discretizations of 1D problems, the number of outliers only depends on the spline order, differential operator and type of boundary conditions [12, 16]. Consequently, the number of eigenvalues computed scales linearly with the dimension and does not depend on the mesh size. For instance, based on the upper bounds provided in [16], for a uniform $C^{p-1}$ discretization of the Laplace on the hypercube $(0,1)^{d}$ with homogeneous Dirichlet boundary conditions, only $d(p-1)$ eigenpairs are required to remove $O(n^{d-1}(p-1))$ outliers, where $n$ is the dimension of the univariate spline space [12]. This method also offers some advantages over the ones presented in [12, 16]. Firstly, its implementation is straightforward: it does not require a change of basis, the standard B-spline basis is sufficient. Secondly, it seems more robust than the method presented in [12], which generally stumbles on domains with curved boundaries, even if the mass matrix is a Kronecker product.
@@ 5 Numerical experiments
This section gathers a few numerical experiments designed to verify our theoretical results and demonstrate the usefulness of our strategies in the context of explicit dynamics. All experiments in this section are done using GeoPDEs [61], a software package for isogeometric analysis.
@@ 5.1 Single-patch geometries
We consider a cubic discretization of the 2D Laplace on two nontrivial single-patch domains shown in Figures 0(a) and 1(a): a stretched square and a quarter of a plate with a hole, represented by a (near) singular NURBS patch. Here, $\mathcal{M}\in\mathcal{S}_{(n_{1},n_{2})}^{+}$ and we construct the first three lumped mass matrices of the sequence. The spectrum of $(\mathcal{K},\mathcal{M})$ and $(\mathcal{K},\mathcal{P}_{i})$, for $i=1,2,3$, is shown in Figures 0(b) and 1(b) for the stretched square and the plate with a hole, respectively. As predicted, the generalized eigenvalues of the matrix pairs $(\mathcal{K},\mathcal{P}_{i})$ monotonically converge to the eigenvalues of $(\mathcal{K},\mathcal{M})$ from below for increasing values of $i$. This property holds for all eigenvalues, including the “outliers”, now characterized by a sharp but smooth increase of the spectrum rather than a stepwise increase. For this reason, the distinction between “outlier” and “regular” eigenvalues is rather ambiguous. For nontrivial problems, “outlier” eigenvalues are merely large eigenvalues, which can be removed using deflation techniques such as those presented in Section 4.1. In order to assess the practical gains of the procedure, we solve the wave equation on the plate geometry shown in Figure 1(a) over the time span $T=[0,6]$ with the manufactured solution $u(x,y,t)=xy(x+4)(y-4)(x^{2}+y^{2}-1)(2+\sin(2\pi t))$. The numerical solutions are computed with the central difference method using the critical time step (1.1) multiplied by a safeguarding factor of $0.85$. The results are shown in Figure 5.3 for a small time ($t=0.65$) and a larger time ($t=2.65$). Figure 5.4 represents the evolution of the $L^{2}$ error over time. As one could expect, increasing the block bandwidth improves the accuracy of the lumping techniques.
Figure 5.5 shows the ratio $(N_{s}+N_{i})/N_{w}$ for the block diagonal matrix $\mathcal{P}_{1}$, where $N_{s}$ and $N_{w}$ are the number of iterations with and without scaling, respectively, and $N_{i}$ is the number of iterations needed by the eigensolver for computing the scaling. The ratio is computed for ranks $r=10,20,40$ and simulations ranging over increasingly larger time spans. The horizontal line for $r=0$ indicates the absence of scaling and serves as comparison. A ratio strictly larger than $1$ indicates a deficit: the saving due to the scaling could not offset the cost for computing it. This situation commonly occurs for short-time simulations. However, the workload for computing a few eigenpairs is quickly amortized over longer simulations and may eventually save more than $50\%$ iterations. In general, one should favor smaller ranks for shorter simulations and larger ranks for longer simulations. Moreover, we could verify over the range of ranks tested that the scaling did not have any significant effect on the quality of the solution.
@@
We now check the convergence of the smallest eigenfrequency for the approximations introduced in Section 3. We consider two test cases designed to evaluate the effect of a coefficient function and geometry mapping. The first one is the unit square with a non-separable (but continuous) density function $\rho(x)=|\sin(xy)|+x+y+1$ and the second one is the quarter of a plate with a hole (see Figure 1(a)). Figures 5(a) and 5(b) show the relative error $\frac{\omega_{1}-\omega_{h,1}}{\omega_{1}}$ for the first eigenfrequency and a cubic discretization. Due to the lack of closed form solutions, the reference eigenfrequency $\omega_{1}$ is a high order approximation computed with the consistent mass on a very fine mesh. The smallest eigenfrequency of $(\mathcal{K},\mathcal{M})$ converges at the expected rate of $2p$, while the smallest eigenfrequency of $(\mathcal{K},\mathcal{P}_{i})$ converges at a reduced quadratic rate. This observation is in agreement with the well-known fact that the row-sum technique converges at a reduced quadratic rate, independently of the spline order [6].
@@
Now we consider a quadratic discretization of the 3D Laplace on the magnet domain shown in Figure 6(a) and test the hierarchical lumped mass matrices described in Section 3.3. Their sparsity pattern is shown in Figure 5.8 together with the consistent mass for $N=6$ subdivisions in each parametric direction. Hierarchical mass lumping leads to a significant reduction of the bandwidth and number of nonzero entries, which drastically speeds up sparse direct solvers. For assessing the performance of mass lumping in explicit dynamics, we solve a sequence of $1000$ linear systems with the consistent mass $\mathcal{M}$ and hierarchical lumped mass matrices $\mathcal{H}_{k}$ for $k=1,2,3$ on increasingly fine meshes. The solver relies on sparse Cholesky factorizations computed on the reordered matrices using nested dissection. According to Table 5.1, mass lumping may save several orders of magnitude of computing time, even for relatively small systems. We also noticed that the reordering only slightly reduced the number of nonzero entries in the Cholesky factors and was not the main driver for the enhanced performance. Table 5.2 shows the computing time for a simulation spanning $50$ seconds. While Table 5.1 only accounts for the effect of the linear system solver, Table 5.2 additionally accounts for the saving in the number of time steps thanks to the increase of the critical time step computed with (1.1).
The impact of hierarchical mass lumping on the accuracy of the solution is (partly) determined by the generalized eigenpairs of $(\mathcal{K},\mathcal{H}_{k})$. The eigenvalues are shown in Figure 6(b) alongside those of $(\mathcal{K},\mathcal{M})$ for $N=6$ subdivisions. Interestingly, $\mathcal{H}_{2}$ seems much more accurate than $\mathcal{H}_{3}$ and yet barely increases its associated CFL condition. This encouraging result indicates that improved accuracy is possible with only a marginal increase in computational cost. Moreover, similarly to Example 5.3, we verified that hierarchical mass lumping delivered second order convergent eigenvalues.
(a) Geometry. The red dot identifies the measurement point in Figure 5.9
@@
Table 5.1: System size and computing times (in seconds) for solving a sequence of $1000$ linear systems with the consistent mass and hierarchical lumped mass matrices. $N=h^{-1}$ denotes the number of subdivisions in each parametric direction.
@@
Table 5.2: System size, computing times (in seconds) and number of time steps for simulating $50$ seconds with the consistent mass and hierarchical lumped mass matrices. $N=h^{-1}$ denotes the number of subdivisions in each parametric direction.
@@
Figure 5.9: Vertical displacement at the point identified by a red dot in Figure 6(a)
@@ 5.2 Multipatch geometries
We consider a cubic discretization of the Laplace on the quarter of a plate with a hole, as shown in Figure 1(a), and split into two patches to remove sources of singularity. The sparsity patterns of the consistent mass and lumped mass matrices $\mathcal{P}_{i}$ for $i=1,2$ are shown in Figure 5.10 for $15$ subdivisions in each parametric direction and each patch. Mass lumping for multipatch problems does not completely remove the coupling but instead focuses on reducing the bandwidth of the diagonal blocks in order to easily form the Schur complement. As explained in Section 4, we try scaling the local stiffness matrices before assembling them into a global matrix. For clarity, we denote $\bar{\mathcal{K}}_{\textrm{loc}}$ the locally scaled stiffness matrix. Given its heuristic nature, this scaling strategy might effect the smallest eigenvalues. Nevertheless, Figure 10(a), obtained for a rank 20 patchwise scaling, reveals that this effect is very mild. By comparing the results with a rank 40 global mass scaling, we notice that the local scaling method removes fewer outliers. This is expected given that it cannot remove interior outliers, arising from the $C^{0}$ coupling of patch interfaces. The convergence test carried out in Figure 10(b) further indicates that the smallest eigenfrequency converges at a second order rate, as it does in the purely lumped mass case (using $P_{1}$). For the sake of clarity, we have only reported the results for $\mathcal{P}_{2}$ but they seem to hold more generally.
@@
This experiment aims at better understanding how patch configurations effect the scaling outcome. Our benchmarks consist of the unit square split into a $4\times 4$ grid of patches and a rectangle of length $4$ and width $0.25$ split into a $1\times 16$ grid. Both configurations are discretized using quadratic B-splines and $N=8$ subdivisions in each direction and each patch. The size and number of patches of both configurations are the same (16 patches of size $0.25\times 0.25$) but are coupled very differently. Although homogeneous Dirichlet boundary conditions are prescribed along the entire boundary in both cases, the first one features a number of interior patches and consequently more degrees of freedom. Figures 11(a) and 11(b) compare for both configurations a local patchwise scaling of rank $10$ with a global scaling of rank $160$. Surprisingly, our local scaling strategy performs poorly even for the second, weakly coupled, configuration. In both cases, the global scaling is again more efficient at removing outliers. However, this strategy is also more expensive than our local (sequential) patchwise scaling, as shown in Tables 5.3 and 5.4 for increasingly fine meshes. The global scaling strategy is slightly faster on the second configuration, probably owing to the smaller system size. Nevertheless, the outcome generally depends on the patch configuration, which effects the spectral properties of the system matrices and consequently the convergence speed of the Lanczos method. Therefore, a direct comparison of global and local scaling is not trivial, even while neglecting potential for parallelism. A more thorough study is needed before drawing definite conclusions.
@@
Time [s]
$N$ Size Global Local
4 361 0.04 0.03
8 1225 0.22 0.06
12 2601 0.48 0.07
16 4489 0.77 0.12
20 6889 1.40 0.16
Table 5.3: Unit square split into a $4\times 4$ grid of patches.
Time [s]
$N$ Size Global Local
4 316 0.02 0.03
8 1144 0.14 0.09
12 2484 0.38 0.12
16 4336 0.83 0.13
20 6700 1.29 0.17
Table 5.4: Rectangle of length $4$ and width $0.25$ split into a $1\times 16$ grid of patches.
@@
As we have noted in Section 3.4, any suitable mass lumping technique may be applied patchwise before assembling the global multipatch lumped mass matrix. In dimension $d\geq 3$, hierarchical mass lumping stands out as the natural candidate. When the context is clear, we do not distinguish the single-patch matrices from their multipatch counterpart. We solve the Laplace eigenvalue problem on the $3$-patch twisted box geometry shown in Figure 12(a), discretized with quadratic B-splines and $N=6$ subdivisions in each parametric direction and each patch. The spectrum of $(\mathcal{K},\mathcal{M})$ and $(\mathcal{K},\mathcal{H}_{k})$ for $k=1,2,3$ is shown in Figure 12(b) and confirm the improved accuracy on the low-frequency part of the spectrum with respect to the row-sum technique (i.e. $\mathcal{H}_{3}$). Similarly to Example 5.4, hierarchical mass lumping yields a drastic reduction of the bandwidth and number of nonzero entries, as shown in Figure 5.14.

### v2

@@ 4.2 Eigenvalue and eigenvector computations
Algorithm 1: Lanczos method [58, Algorithm 9.1]
The vectors $\mathbf{v}_{j}$ computed during the course of the iterations serve to build a $B$-orthonormal basis for a Krylov subspace from which eigenvalue and eigenvector approximations are extracted. The interested reader may refer to standard textbooks detailing this procedure; e.g. [56, 58].
@@
Although we have implemented an eigensolver from scratch for the sake of writing this paper, it is absolutely not necessary for applying the techniques presented herein. Several efficient implementations are available in software packages and all the user has to worry about is supplying an algorithm for computing matrix-vector multiplications with the (implicitly defined) stiffness matrix and solving linear systems with the (lumped) mass matrix, which depends on the nature of the problem (e.g. single-patch, multipatch,…). These two operations are reminiscent of the central difference method and as a matter of fact, if the number of iterations $m$ remains relatively small, which is ensured through restarting procedures, each iteration of the Lanczos method costs nearly as much as an iteration of the central difference method. Moreover, the Lanczos method is known to converge very fast to eigenvalues that are well separated from the rest of the spectrum [63, 56, 58], which is precisely a distinctive feature of outliers. However, the time span of the simulation must be sufficiently large to amortize the cost of computing outlier eigenpairs. In general, the shorter the simulation, the smaller the number of outliers we can afford computing. In practice, their number is further limited by computer resources. For very large problems, simply storing all outlier eigenvectors is plainly infeasible. Following common guidelines relating the size of the Krylov subspace to the number of eigenpairs [61], if one can afford storing $2k$ vectors, then one might compute roughly $k$ eigenpairs. Finally, although our method scales down outlier frequencies, it does not remove the corresponding outlier modes that might negatively impact the solution.
The deflation procedure proposed in this section is very general and can in principle be applied to any type of problem, including nontrivial (multipatch) geometries. In the multipatch setting, it is possible to locally scale the single-patch system matrices, similarly to local (elementwise) mass scaling techniques [27, 64]. Indeed, Lemma 3.23 reveals that the largest eigenvalues of $(\mathcal{K},\mathcal{P}_{i})$ could be controlled by the largest eigenvalues of $(\mathcal{K}_{r},\mathcal{P}_{r,i})$ and suggests a scaling strategy directly targeting the origin of the issue: at the patch level. While this strategy is generally cheaper than scaling $(\mathcal{K},\mathcal{P}_{i})$ globally, it has three shortcomings: firstly, the assembly into global matrices generally affects the smallest eigenvalues; secondly it cannot remove outliers introduced by the $C^{0}$ coupling of patch interfaces and finally, since the inverse of the global mass matrix is not given by the assembly of the local inverses, we cannot directly apply (4.1) locally. To resolve the third issue, we suggest locally scaling the stiffness matrix instead. Denoting $\bar{\mathcal{K}}_{r}$ the locally scaled stiffness matrix and recalling that the perturbation is negative semidefinite (see Section 4.1), $\bar{\mathcal{K}}_{r}\preceq\mathcal{K}_{r}$ and
@@
For some special cases it is yet far more advantageous to exploit the structure of the problem. In particular, for problems featuring a Kronecker product structure, all outliers can be removed by separately scaling the 1D factor matrices. For maximally smooth spline discretizations of 1D problems, the number of outliers only depends on the spline order, differential operator and type of boundary conditions [12, 16]. Consequently, the number of eigenvalues computed scales linearly with the dimension and does not depend on the mesh size. For instance, based on the upper bounds provided in [16], for a uniform $C^{p-1}$ discretization of the Laplace on the hypercube $(0,1)^{d}$ with homogeneous Dirichlet boundary conditions, at most $dp$ eigenpairs are required to remove $O(n^{d-1}(p-1))$ outliers, where $n$ is the dimension of the univariate spline space [12]. This method also offers some advantages over the ones presented in [12, 16]. Firstly, its implementation is straightforward: it does not require a change of basis, the standard B-spline basis is sufficient. Secondly, our strategy only relies on algebraic properties and not geometric ones, contrary to the method presented in [12], which generally stumbles on domains with curved boundaries, even if the mass matrix is a Kronecker product.
@@ 5 Numerical experiments
This section gathers a few numerical experiments designed to verify our theoretical results and demonstrate the usefulness of our strategies in the context of explicit dynamics. All experiments in this section are done using GeoPDEs [65], an open source Matlab/Octave software package for isogeometric analysis.
@@ 5.1 Single-patch geometries
We consider a cubic discretization of the 2D Laplace on two nontrivial single-patch domains shown in Figures 1(a) and 2(a): a stretched square and a quarter of a plate with a hole, represented by a (near) singular NURBS patch. These geometries are discretized with $(20,20)$ and $(40,20)$ subdivisions, respectively. Here, $\mathcal{M}\in\mathcal{S}_{(n_{1},n_{2})}^{+}$ and we construct the block lumped matrices $\mathcal{P}_{i}$ for $i=1,2,3$. The spectrum of $(\mathcal{K},\mathcal{M})$ and $(\mathcal{K},\mathcal{P}_{i})$, for $i=1,2,3$, is shown in Figures 1(b) and 2(b) for the stretched square and the plate with a hole, respectively. As predicted, the generalized eigenvalues of the matrix pairs $(\mathcal{K},\mathcal{P}_{i})$ monotonically converge to the eigenvalues of $(\mathcal{K},\mathcal{M})$ from below for increasing values of $i$. This property holds for all eigenvalues, including the “outliers”, now characterized by a sharp but smooth increase of the spectrum rather than a stepwise increase. For this reason, the distinction between “outlier” and “regular” eigenvalues is rather ambiguous. For nontrivial problems, “outlier” eigenvalues are merely large eigenvalues, which can be removed using deflation techniques such as those presented in Section 4.1. In order to assess the practical gains of the procedure, we solve the wave equation on the plate geometry shown in Figure 2(a) over the time span $T=[0,6]$ with the manufactured solution $u(x,y,t)=xy(x+4)(y-4)(x^{2}+y^{2}-1)(2+\sin(2\pi t))$. The numerical solutions are computed with the central difference method using the critical time step (1.1) multiplied by a safeguarding factor of $0.85$. The results are shown in Figure 5.3 for a small time ($t=0.65$) and a larger time ($t=2.65$). Figure 5.4 represents the evolution of the $L^{2}$ error over time. As one could expect, increasing the block bandwidth improves the accuracy of the lumping techniques. Unfortunately, the improvement also holds true for the outliers, which must be removed. In order to evaluate the computational savings of the outlier removal technique, we consider the ratio $(N_{s}+N_{i})/N_{w}$ for fixed rank values $r$, where $N_{s}$ and $N_{w}$ are the number of iterations with and without scaling, respectively, and $N_{i}$ is the number of iterations needed by the eigensolver for computing the $r+1$ largest eigenpairs. As we have seen in Section 4.2, an iteration of Lanczos costs nearly as much as an iteration of the central difference method and justifies adding $N_{i}$ to $N_{s}$. For conciseness, Figure 5.5 only reports the results for the block diagonal matrix $\mathcal{P}_{1}$ and rank values $r=10,20,40$ for a convergence tolerance of $10^{-3}$. We could verify over the range of ranks tested that the scaling did not have any significant effect on the quality of the solution. The horizontal line for $r=0$ in Figure 5.5 indicates the absence of scaling and serves as comparison. A ratio strictly larger than $1$ indicates a deficit: the saving due to the scaling could not offset the cost for computing it. This situation commonly occurs for short-time simulations. However, the workload for computing a few eigenpairs is quickly amortized over longer simulations and may eventually save more than $50\%$ iterations. In general, one should favor smaller ranks for shorter simulations and larger ranks for longer simulations. In practice, fixing the total number of Lanczos iterations and a loose target rank (defining the size of the Krylov subspace) should help keep control of the computational overhead.
@@
We now check the convergence of the smallest eigenfrequency for the approximations introduced in Section 3. We consider two test cases designed to evaluate the effect of a coefficient function and geometry mapping. The first one is the unit square with a non-separable (but continuous) density function $\rho(x,y)=|\sin(xy)|+x+y+1$ and the second one is the quarter of a plate with a hole (see Figure 2(a)). Figures 6(a) and 6(b) show the relative error $\frac{\omega_{1}-\omega_{h,1}}{\omega_{1}}$ for the first eigenfrequency and a cubic discretization. Due to the lack of closed form solutions, the reference eigenfrequency $\omega_{1}$ is a high order approximation computed with the consistent mass on a very fine mesh. The smallest eigenfrequency of $(\mathcal{K},\mathcal{M})$ converges at the expected rate of $2p$, while the smallest eigenfrequency of $(\mathcal{K},\mathcal{P}_{i})$ converges at a reduced quadratic rate. This observation is in agreement with the well-known fact that the row-sum technique converges at a reduced quadratic rate, independently of the spline order [6].
@@
Now we consider a quadratic discretization of the 3D Laplace on the magnet domain shown in Figure 7(a) and test the hierarchical lumped mass matrices described in Section 3.3. Their sparsity pattern is shown in Figure 5.8 together with the consistent mass for $N=12$ subdivisions in each parametric direction. Hierarchical mass lumping leads to a significant reduction of the bandwidth and number of nonzero entries, which drastically speeds up sparse direct solvers. For assessing the performance of mass lumping in explicit dynamics, we solve a sequence of $1000$ linear systems with the consistent mass $\mathcal{M}$ and hierarchical lumped mass matrices $\mathcal{H}_{k}$ for $k=1,2,3$ on increasingly fine meshes. The solver relies on sparse Cholesky factorizations computed on the reordered matrices using nested dissection. According to Table 1(a), mass lumping may save several orders of magnitude of computing time, even for relatively small systems. We also noticed that the reordering only slightly reduced the number of nonzero entries in the Cholesky factors and was not the main driver for the enhanced performance. Table 1(b) shows the computing time for a simulation spanning $50$ seconds. While Table 1(a) only accounts for the effect of the linear system solver, Table 1(b) additionally accounts for the saving in the number of time steps thanks to the increase of the critical time step computed with (1.1).
The impact of hierarchical mass lumping on the accuracy of the solution is (partly) determined by the generalized eigenpairs of $(\mathcal{K},\mathcal{H}_{k})$. The eigenvalues are shown in Figure 7(b) alongside those of $(\mathcal{K},\mathcal{M})$ for $N=12$ subdivisions. Interestingly, $\mathcal{H}_{2}$ seems much more accurate than $\mathcal{H}_{3}$ and yet barely increases its associated CFL condition. This encouraging result indicates that improved accuracy is possible with only a marginal increase in computational cost. Moreover, similarly to Example 5.3, we verified that hierarchical mass lumping delivered second order convergent eigenvalues.
(a) Geometry
@@
(a) Sequence of $1000$ linear systems
@@
(b) Simulation of $50$ seconds
Table 5.1: System size, computing times (in seconds) and number of time steps for the consistent mass and hierarchical lumped mass matrices. $N=h^{-1}$ denotes the number of subdivisions in each parametric direction.
@@
Figure 5.9: $L^{2}$ error with respect to the solution for the consistent mass
@@ 5.2 Multipatch geometries
We consider a cubic discretization of the Laplace on the quarter of a plate with a hole, as shown in Figure 2(a), and split into two patches to remove sources of singularity. The sparsity patterns of the consistent mass and lumped mass matrices $\mathcal{P}_{i}$ for $i=1,2$ are shown in Figure 5.10 for $15$ subdivisions in each parametric direction and each patch. Mass lumping for multipatch problems does not completely remove the coupling but instead focuses on reducing the bandwidth of the diagonal blocks in order to easily form the Schur complement. While increasing the bandwidth generally improves the accuracy, forming the Schur complement also typically takes a heavier toll. This tradeoff must be carefully balanced on a case-by-case basis. As explained in Section 4, we try scaling the local stiffness matrices before assembling them into a global matrix. For clarity, we denote $\bar{\mathcal{K}}_{\textrm{loc}}$ the locally scaled stiffness matrix. Given its heuristic nature, this scaling strategy might affect the smallest eigenvalues. Nevertheless, Figure 11(a), obtained for a rank 20 patchwise scaling and a tolerance of $10^{-3}$, reveals that this effect is very mild. By comparing the results with a rank 40 global mass scaling, we notice that the local scaling method removes fewer outliers. This is expected given that it cannot remove interior outliers, arising from the $C^{0}$ coupling of patch interfaces. The convergence test carried out in Figure 11(b) further indicates that the smallest eigenfrequency converges at a second order rate, as it does in the purely lumped mass case (using $P_{1}$). For the sake of clarity, we have only reported the results for $\mathcal{P}_{2}$ but they seem to hold more generally.
@@
This experiment aims at better understanding how patch configurations affect the scaling outcome. Our benchmarks consist of the unit square split into a $4\times 4$ grid of patches and a rectangle of length $4$ and width $0.25$ split into a $1\times 16$ grid. Both configurations are discretized using quadratic B-splines and $N=8$ subdivisions in each direction and each patch. The size and number of patches of both configurations are the same (16 patches of size $0.25\times 0.25$) but are coupled very differently. Although homogeneous Dirichlet boundary conditions are prescribed along the entire boundary in both cases, the first one features a number of interior patches and consequently more degrees of freedom. Figures 12(a) and 12(b) compare for both configurations a local patchwise scaling of rank $10$ with a global scaling of rank $160$. Surprisingly, our local scaling strategy performs poorly even for the second, weakly coupled, configuration. In both cases, the global scaling is again more efficient at removing outliers. However, this strategy is also more expensive than our local (sequential) patchwise scaling, as shown in Tables 2(a) and 2(b) for increasingly fine meshes. The global scaling strategy is generally slightly faster on the second configuration, probably owing to the smaller system size. Nevertheless, the outcome generally depends on the patch configuration, which affects the spectral properties of the system matrices and consequently the convergence speed of the Lanczos method. Therefore, a direct comparison of global and local scaling is not trivial, even while neglecting potential for parallelism. A more thorough study is needed before drawing definite conclusions. In the standard tensor product case, mesh refinement and degree elevation widens the gap between “outlier” and “regular” eigenvalues and benefits Lanczos while in these multipatch examples, the number of Lanczos iterations (for a tolerance of $10^{-3}$) remains roughly constant beyond a certain system size and may occasionally exceed the system size (especially for small systems) due to restarting.
@@
Time [s] Iterations
$N$ Size Global Local Global Local
4 361 0.05 0.03 480 160
8 1225 0.18 0.03 832 80
12 2601 0.45 0.04 800 112
16 4489 0.76 0.09 800 228
20 6889 1.21 0.12 832 145
(a) Unit square split into a $4\times 4$ grid of patches.
Time [s] Iterations
$N$ Size Global Local Global Local
4 316 0.02 0.02 316 157
8 1144 0.10 0.04 480 81
12 2484 0.37 0.05 960 112
16 4336 0.67 0.09 1120 236
20 6700 1.29 0.11 1088 147
(b) Rectangle of length $4$ and width $0.25$ split into a $1\times 16$ grid of patches.
Table 5.2: System size, computing times (in seconds) and number of Lanczos iterations for the global and local (sequential) scaling strategies. Computing times are averaged out over $5$ independent runs. Local iteration counts are averaged out across patches and rounded to the nearest integer. $N=h^{-1}$ denotes the number of subdivisions in each parametric direction and each patch.
@@
As we have noted in Section 3.4, any suitable mass lumping technique may be applied patchwise before assembling the global multipatch lumped mass matrix. In dimension $d\geq 3$, hierarchical mass lumping stands out as the natural candidate. When the context is clear, we do not distinguish the single-patch matrices from their multipatch counterpart. We solve the Laplace eigenvalue problem on the $3$-patch twisted box geometry shown in Figure 13(a), discretized with quadratic B-splines and $N=12$ subdivisions in each parametric direction and each patch. The spectra of $(\mathcal{K},\mathcal{M})$ and $(\mathcal{K},\mathcal{H}_{k})$ for $k=1,2,3$ are shown in Figure 13(b) and confirm the improved accuracy on the low-frequency part of the spectrum with respect to the row-sum technique (i.e. $\mathcal{H}_{3}$). Similarly to Example 5.4, hierarchical mass lumping yields a drastic reduction of the bandwidth and number of nonzero entries, as shown in Figure 5.14.